# Overcoming the Stability Gap in Continual Learning

### v3

@@ Appendix D Additional CIL Analysis & Experiments
In this section, we conduct additional analysis of the CIL experiments in the main results as well as present additional experiments.
@@ D.1 Qualitative Analysis
Stability gap over all rehearsal sessions. In our main text, our figures are averaged across rehearsal sessions. In Fig. 4, we instead present all of the learning curves in sequence, where we denote when the next batch containing new classes is received.
@@
Computational efficiency. To measure computational efficiency, we consider two criteria: 1) number of network updates (Sec. 3.1) and 2) FLOPs (floating-point operations). For FLOPs analysis, we use DeepSpeed  [^5] https://github.com/microsoft/DeepSpeed with the same GPU across compared models. We perform computational analysis for CIL on a combination of ImageNet-1K and Places365-LT datasets.
@@
Figure 5: Computational efficiency. Our method, SGM, provides a 16.7×16.7$\times$16.7 × speedup in number of network updates and a 31.9×31.9$\times$31.9 × speedup in TFLOPs compared to a joint model (upper bound) with the combined 1365 class dataset (ImageNet-1K and Places365-LT combined). For SGM and conventional rehearsal, we show the stability gap in the learning curve averaged over rehearsal sessions.
Stability-plasticity. In Fig. 6, we also illustrate the model’s accuracy on new, old, and all classes in all rehearsal sessions where SGM achieves higher accuracy than vanilla. This indicates that SGM consistently improves the model’s plasticity (Fig. 6(a)), stability (Fig. 6(b)), and knowledge accumulation (Fig. 6(c)).
@@ D.5 Balanced (Non-LT) Dataset
In a real-world setting, data distribution is commonly long-tailed and imbalanced, hence we used Places365-LT dataset in the main results. However, our analysis holds for balanced and non-LT dataset as well. We study this using Places365-Standard. We omit storage constraints in this experiment to solely focus on mitigation methods without the influence of other variables. Results in Table 10 show that SGM outperforms vanilla rehearsal in all criteria.
@@
Figure 7: Comparison with non-rehearsal method. The Y-axis shows an average accuracy of 6 runs with a standard deviation (shaded region). The network is trained on ImageNet-1K and then learns 365 new classes from Places-LT over five batches ($73$ new classes and 600 iterations per batch). When a new batch arrives, accuracy on ImageNet-1k for LwF plummets. LwF fails to recover performance and ends up with a large stability gap. In contrast, LwF with SGM does not plummet like LwF and shows better performance throughout the CL phase with a significantly reduced stability gap.
Appendix E SGM Enhances Non-rehearsal Methods
We conducted an experiment to compare vanilla LwF with a version of LwF that uses SGM without rehearsal during CIL of ImageNet-1K and Places365-LT. Overall results are given in Table 11 and a learning curve is given in Fig. 7. As expected based on prior results, rehearsal methods vastly outperform LwF; however, we find that SGM provides an enormous benefit to LwF in terms of reducing the stability gap, resulting in increased accuracy.
Figure 8: Need for proposed metrics. We present a series of synthetic plots that display the performance of a model on an old task as it learns a new task. The largest drop in accuracy on the old task is marked by a dotted line with an arrow. This drop in accuracy is also displayed as a number for ease of comparison. In sub-figures (a) and (b), our metric delivers a different outcome from theirs. In sub-figures (c) and (d), unlike their metric, our metric can distinguish between two CL models.
@@ Appendix F Reasons for New Metrics
Case 1. The metric proposed in earlier work (De Lange et al., 2023) measures the stability gap of a model by evaluating the largest drop in old task performance compared to the same model’s performance. According to their metric, CL model 1 is considered to be better than CL model 2 as it exhibits a smaller drop in accuracy for the old task, as shown in Fig. 8(a). However, in contrast to this observation, it can be observed that CL model 2 actually performs better than CL model 1 as it shows a smaller drop in old task accuracy compared to CL model 1 while employing the joint model (upper bound) as a universal reference point. This is captured by our metric and can be seen in Fig. 8(b).

### v4

@@ Appendix D Additional CIL Analysis & Experiments
In this section, we conduct additional analysis of the CIL experiments in the main results as well as present additional experiments. Implementation details and dataset details are given in Appendix C and B.
@@ D.1 Qualitative Analysis
D.1.1 Stability Gap Over All Rehearsal Sessions
In our main text, our figures are averaged across rehearsal sessions. In Fig. 4, we instead present all of the learning curves in sequence, where we denote when the next batch containing new classes is received.
@@ D.1.1 Stability Gap Over All Rehearsal Sessions
D.1.2 Computational Efficiency
To measure computational efficiency, we consider two criteria: 1) number of network updates (Sec. 3.1) and 2) FLOPs (floating-point operations). For FLOPs analysis, we use DeepSpeed  [^6] https://github.com/microsoft/DeepSpeed with the same GPU across compared models. We perform computational analysis for CIL on a combination of ImageNet-1K and Places365-LT datasets.
@@ D.1.2 Computational Efficiency
Figure 5: Computational efficiency. Our method, SGM, provides a 16.7×16.7$\times$16.7 × speedup in the number of network updates and a 31.9×31.9$\times$31.9 × speedup in TFLOPs compared to a joint model (upper bound) with the combined 1365 class dataset (ImageNet-1K and Places365-LT combined). For SGM and conventional rehearsal, we show the stability gap in the learning curve averaged over rehearsal sessions.
@@
D.1.3 Stability-Plasticity
In Fig. 6, we also illustrate the model’s accuracy on new, old, and all classes in all rehearsal sessions where SGM achieves higher accuracy than vanilla. This indicates that SGM consistently improves the model’s plasticity (Fig. 6(a)), stability (Fig. 6(b)), and knowledge accumulation (Fig. 6(c)).
@@ D.5 Balanced (Non-LT) Dataset
In a real-world setting, data distribution is commonly long-tailed and imbalanced, hence we used Places365-LT dataset in the main results. However, our analysis holds for balanced and non-LT datasets as well. We study this using Places365-Standard. We omit storage constraints in this experiment to solely focus on mitigation methods without the influence of other variables. Results in Table 10 show that SGM outperforms vanilla rehearsal in all criteria.
@@
Figure 7: Comparison with non-rehearsal method. The Y-axis shows an average accuracy (%) of 6 runs with a standard deviation (shaded region). The network is trained on ImageNet-1K and then learns 365 new classes from Places-LT over five batches (73 new classes and 600 iterations per batch). When a new batch arrives, accuracy on ImageNet-1K for LwF plummets. LwF fails to recover performance and ends up with a large stability gap. In contrast, LwF with SGM does not plummet like LwF and shows better performance throughout the CL phase with a significantly reduced stability gap.
Appendix E SGM Enhances Non-rehearsal Methods
We conducted an experiment to compare vanilla LwF with a version of LwF that uses SGM without rehearsal during CIL of ImageNet-1K and Places365-LT. For this, we use ConvNeXtV2-Femto pre-trained on ImageNet-1K using SSL. After being pre-trained on ImageNet-1K, the model sequentially learns 5 incremental batches of data (73 classes per batch) from Places365-LT in offline CL with CIL ordering. The training is compute-constrained where the model is updated over 600 minibatches. Each minibatch consists of 64 new samples and rehearsal of old samples is omitted. Additional implementation details are given in Appendix C.
Overall results are given in Table 11 and a learning curve is given in Fig. 7. As expected based on prior results, rehearsal methods vastly outperform LwF; however, we find that SGM provides an enormous benefit to LwF in terms of reducing the stability gap, resulting in increased accuracy.
@@ Appendix F Reasons for New Metrics
Case 1. The metric proposed in earlier work (De Lange et al., 2023) measures the stability gap of a model by evaluating the largest drop in old task performance compared to the same model’s performance. According to their metric, CL model 1 is considered to be better than CL model 2 as it exhibits a smaller drop in accuracy for the old task, as shown in Fig. 8(a). However, in contrast to this observation, it can be observed that CL model 2 performs better than CL model 1 as it shows a smaller drop in old task accuracy compared to CL model 1 while employing the joint model (upper bound) as a universal reference point. This is captured by our metric and can be seen in Fig. 8(b).
@@
Figure 8: Efficacy of our metric. We present a series of synthetic plots that display the performance of a model on an old task as it learns a new task. The largest drop in accuracy on the old task is marked by a dotted line with an arrow. This drop in accuracy is also displayed as a number for ease of comparison. In sub-figures (a) and (b), our metric delivers a different outcome from theirs (prior work). In sub-figures (c) and (d), unlike their metric, our metric can distinguish between two CL models.
Appendix G Soft Targets Update
Figure 9: Process of updating soft targets. Here we illustrate the update mechanism of our dynamic soft targets. For each class $k$, we maintain a probability distribution over total $K$ classes, denoted by $\mathbf{u}_{k}\in\mathbb{R}^{K}$. During iteration $i$, we update this vector $\mathbf{u}_{k}^{i}$ if the model ($\theta_{i}$) correctly predicts the class $k$. In the following iteration $i+1$, $\mathbf{u}_{k}^{i}$ will serve as soft targets to train the model, $\theta_{i+1}$. Unlike hard targets, soft targets do not enforce strict inter-class independence and prevent the stability gap.
In the main text in Sec. 4, we proposed dynamic soft targets to prevent the stability gap. In this section, we illustrate how we update the soft targets in Fig. 9. Soft targets are updated using the model’s predicted probabilities.
Appendix H Additional Discussion
What is the connection between evaluation metrics and stability gap mitigation approaches?
In the main text in Sec. 3.2, we formalized our evaluation metrics to measure the stability gap ($\mathcal{S}_{\Delta}$), plasticity gap ($\mathcal{P}_{\Delta}$), and continual knowledge gap ($\mathcal{CK}_{\Delta}$). We found that our proposed mitigation methods in Sec. 4 effectively mitigated the stability, plasticity, and continual knowledge gaps when evaluated using metrics from Sec. 3.2. Therefore, the evaluation metrics can be regarded as the tools to measure the efficacy of the proposed mitigation methods.
Mitigation methods were designed to prevent factors contributing to the stability gap such as large loss at the output layer and excessive network plasticity. All three metrics were used in all evaluations and experiments in the main paper and the Appendix.